# Dimitrios Kostas [^1] Mathematics subject classification:65M60,65M12,35K55

### v2

@@ 4.2 Assumptions for All Numerical Experiments
For time discretization, we employ the backward Euler method, which is first-order accurate and unconditionally stable, making it well-suited for stiff, time-dependent problems. To handle the nonlinear terms, we apply Picard linearization—a fixed-point iterative scheme that avoids derivative computations required in Newton methods. Although Picard iteration may converge more slowly than Newton’s method, it provides sufficient accuracy for our problem, particularly when initialized close to the true solution.
@@ 4.2.1 Numerical examples
Table 7: observed norm values indicate a trend towords instability the norms $\|u_{h1}\|^{2}_{L^{2}(L^{2}(\Omega))}$ and $\|u_{h2}\|^{2}_{L^{2}(L^{2}(\Omega))}$ grow significantly. Despite their moderate magnitude, these increasing values suggest that the discrete solution’s norms are not uniformly bounded. Therefore, while the scheme remains in an intermediate regime—not definitively stable nor outright unstable—the results highlight that stability is sensitive to the choice of penalty parameter in conjunction with $a$ and $b$. Next setting the coefficients $b=0.8$ and the reaction diffusion constant $a=0.8$ when penalty $penal=10^{8}$
@@
Table 11: From our examination of the matrices associated with the Nonsymmetric Interior Penalty Galerkin (NIPG) and Symmetric Interior Penalty Galerkin (SIPG) methods, we observe that the stability results coincide under the conditions considered. This agreement is not inherent to the methods alone; rather, it also depends critically on the reaction–diffusion parameters, the coefficients governing the non linearity, and—importantly—the penalty parameter. In our study, we employed an exceedingly large penalty parameter of the order $10^{8}$. As a result, both NIPG and SIPG methods exhibit equivalent stability behavior within our computational framework.
The case of Incomplete Interrior Penalty Galerkin, Now we extend the comparison between the Symmetric Interior Penalty Galerkin (SIPG) and Nonsymmetric Interior Penalty Galerkin (NIPG) methods by including the Incomplete Interior Penalty Galerkin (IIPG) variant.For all three methods, we investigate their behavior in the regime of exceedingly large penalty parameters. Specifically, by selecting a penalty of the order $10^{8}$, we assess and compare the previous resulting stability properties with Incomplete Interrior Penalty Galerkin (IIPG) formulations.
@@ 5.1 Discussion
However, it is important to note that despite the stabilizing influence of the penalty term, instability was still observed in certain cases. These instabilities appear to result from specific combinations of the reaction–diffusion coefficients and the nonlinear term’s strength, highlighting the delicate interplay between the various parameters in the system.
@@ 5.3 Future Work and Open Problems
  • Is the resulting space–time DG scheme stable when applied to the complex Landau equation?
@@
These questions are not only mathematically rich but also practically significant. For example, if the computational domain is altered—from a simple interval to a more intricate or irregular geometry—will the numerical method still retain its stability and convergence properties? Understanding this behavior is essential for realistic applications where domain boundaries may evolve over time or exhibit structural complexity. From the broader perspective of numerical analysis, future investigations should aim to analyze these methods under minimal regularity assumptions, especially in cases where the solution exhibits low smoothness or singularities. These are particularly challenging regimes where classical numerical techniques often falter. In this setting, emerging tools from Artificial Intelligence (AI) offer promising avenues for advancement. For instance:
@@
In summary, this work not only introduces a rigorous and flexible DG-based numerical framework for solving the complex Landau equation but also lays the groundwork for addressing a wide range of open theoretical and computational problems. These challenges form the basis for future research directions in numerical analysis, computational mathematics, and scientific computing—fields where innovation in methodology and theory continues to be essential.

### v3

@@ 4.2 Assumptions for All Numerical Experiments
For time discretization, we employ the backward Euler method, which is first-order accurate and unconditionally stable, making it well-suited for stiff, time-dependent problems. To handle the nonlinear terms, we apply Picard linearization a fixed-point iterative scheme that avoids derivative computations required in Newton methods. Although Picard iteration may converge more slowly than Newton’s method, it provides sufficient accuracy for our problem, particularly when initialized close to the true solution.
@@ 4.2.1 Numerical examples
Table 7: observed norm values indicate a trend towords instability the norms $\|u_{h1}\|^{2}_{L^{2}(L^{2}(\Omega))}$ and $\|u_{h2}\|^{2}_{L^{2}(L^{2}(\Omega))}$ grow significantly. Despite their moderate magnitude, these increasing values suggest that the discrete solution norms are not uniformly bounded. Therefore, while the scheme remains in an intermediate regime not definitively stable nor outright unstable ,the results highlight that stability is sensitive to the choice of penalty parameter in conjunction with $a$ and $b$. Next setting the coefficients $b=0.8$ and the reaction diffusion constant $a=0.8$ when penalty $penal=10^{8}$
@@
Table 11: From our examination of the matrices associated with the Nonsymmetric Interior Penalty Galerkin (NIPG) and Symmetric Interior Penalty Galerkin (SIPG) methods, we observe that the stability results coincide under the conditions considered. This agreement is not inherent to the methods alone; rather, it also depends critically on the reaction diffusion parameters, the coefficients governing the non linearity, and importantly the penalty parameter. In our study, we employed an exceedingly large penalty parameter of the order $10^{8}$. As a result, both NIPG and SIPG methods exhibit equivalent stability behavior within our computational framework.
The case of Incomplete Interior Penalty Galerkin, Now we extend the comparison between the Symmetric Interior Penalty Galerkin (SIPG) and Nonsymmetric Interior Penalty Galerkin (NIPG) methods by including the Incomplete Interior Penalty Galerkin (IIPG) variant.For all three methods, we investigate their behavior in the regime of exceedingly large penalty parameters. Specifically, by selecting a penalty of the order $10^{8}$, we assess and compare the previous resulting stability properties with Incomplete Interior Penalty Galerkin (IIPG) formulations.
@@ 5.1 Discussion
However, it is important to note that despite the stabilizing influence of the penalty term, instability was still observed in certain cases. These instabilities appear to result from specific combinations of the reaction-diffusion coefficients and the nonlinear term’s strength, highlighting the delicate interplay between the various parameters in the system.
@@ 5.3 Future Work and Open Problems
  • Is the resulting space-time DG scheme stable when applied to the complex Landau equation?
@@
These questions are not only mathematically rich but also practically significant. For example, if the computational domain is altered from a simple interval to a more intricate or irregular geometry will the numerical method still retain its stability and convergence properties? Understanding this behavior is essential for realistic applications where domain boundaries may evolve over time or exhibit structural complexity. From the broader perspective of numerical analysis, future investigations should aim to analyze these methods under minimal regularity assumptions, especially in cases where the solution exhibits low smoothness or singularities. These are particularly challenging regimes where classical numerical techniques often falter. In this setting, emerging tools from Artificial Intelligence (AI) offer promising avenues for advancement. For instance:
@@
In summary, this work not only introduces a rigorous and flexible DG-based numerical framework for solving the complex Landau equation but also lays the groundwork for addressing a wide range of open theoretical and computational problems. These challenges form the basis for future research directions in numerical analysis, computational mathematics, and scientific computing-fields where innovation in methodology and theory continues to be essential.